An infinite, nonconducting sheet has a surface charge density σ = +7.18 pC/m2. (a) How much work is done by the electric field due to the sheet if a particle of charge q0 = 1.60 × 10-19 C is moved from the sheet to a point P at distance d = 2.40 cm from the sheet? (b) If the electric potential V is defined to be zero on the sheet, what is V at P?
Answer #1

The workdone by the electric field is given by

W =Fx = qEx =\frac{q\sigma x}{\epsilon_0}=\frac{1.6*10^{-19}*7.18*10^{-12}*0.024}{8.854*10^{-12}}=19.46meV\\

The potential at point P is

V = W/q =W/e=19.46mV\\